Solve by the substitution method 3x+7y=40
-4x+y=19

Answer by LinnW(1048)

3x+7y=40
-4x+y=19
Solve for y in the second equation
-4x+y=19
add 4x to each side
y = 4x + 19
Now substitute (4x + 19) for y in 3x+7y=40
3x + 7(4x + 19) = 40
3x + 28x + 133 = 40
31x + 133 = 40
add -133 to each side
31x = -93
divide by 31
x = -93/31
x = -3
Since y = 4x + 19 ,
y = 4(-3) + 19
y = -12 + 19
y = 7